OpenStudy (anonymous):

word problems need help badly what is the difference btw these two tempuratures sahara desert one day it was 136 f and gobi desert is -50 f was recorded ?

OpenStudy (curry):

186

OpenStudy (anonymous):

danm u good

OpenStudy (anonymous):

lol I have more

OpenStudy (anonymous):

i just want to get this over and done with stupid math packet

OpenStudy (anonymous):

136 - (-50) = 136 + 50 = 186

OpenStudy (anonymous):

thanks canu help me

OpenStudy (anonymous):

I have more word problems

OpenStudy (anonymous):

1. the punic wars began in 264 b.c. and ended in 146 b.b how long did the punic wars last ?

OpenStudy (anonymous):

2. Roman Civil began in 509 b.c and ended in 476 a.d how long did Roman Civ last ?

OpenStudy (anonymous):

3. In buffo new york the temp was 14 f in the morning. if the temp dropped 7 f what is the temp now?